In a laboratory, there were just enough microscopes, test tubes, and calculators so that every 3 students had to share a microscope, every 4 students had to share a test tube, and every 5 students had to share a calculator. If the sum of the number of microscopes, test tubes, and calculators used by the class was 94, how many students were in the class?
Answer by stanbon(75887)

In a laboratory, there were just enough microscopes, test tubes, and calculators so that every 3 students had to share a microscope, every 4 students had to share a test tube, and every 5 students had to share a calculator. If the sum of the number of microscopes, test tubes, and calculators used by the class was 94, how many students were in the class?
-----
# of microscopes = x/3
# of test tubes = x/4
# of calculators = x/5
Solve:
x/3 + x/4 + x/5 = 94
--
Multiply thru by 60 to get:
20x + 15x + 12x = 60*94
---
47x = 60*94
x = 60*2
x = 120 (# of students)
------------
Cheers,
Stan H.






Answer: 120 students